# Noise-based reward-modulated learning

### v1

@@ Abstract
Recent advances in reinforcement learning (RL) have led to significant improvements in task performance. However, training neural networks in an RL regime is typically achieved in combination with backpropagation, limiting their applicability in resource-constrained environments or when using non-differentiable neural networks. While noise-based alternatives like reward-modulated Hebbian learning (RMHL) have been proposed, their performance has remained limited, especially in scenarios with delayed rewards, which require retrospective credit assignment over time. Here, we derive a novel noise-based learning rule that addresses these challenges. Our approach combines directional derivative theory with Hebbian-like updates to enable efficient, gradient-free learning in RL. It features stochastic noisy neurons which can approximate gradients, and produces local synaptic updates modulated by a global reward signal. Drawing on concepts from neuroscience, our method uses reward prediction error as its optimization target to generate increasingly advantageous behavior, and incorporates an eligibility trace to facilitate temporal credit assignment in environments with delayed rewards. Its formulation relies on local information alone, making it compatible with implementations in neuromorphic hardware. Experimental validation shows that our approach significantly outperforms RMHL and is competitive with BP-based baselines, highlighting the promise of noise-based, biologically inspired learning for low-power and real-time applications.
@@ 1 Introduction
Noise-based learning methods, which leverage stochastic fluctuations in neural activity [15, 20, 54, 52, 13] or synaptic weights [10, 57, 5, 8] to approximate gradients, offer a pathway toward simpler and more light-weight optimization techniques, that may reflect mechanisms within the brain itself [43, 29]. Given the prevalence of noise in biological neural systems [9, 45, 31], using noise as a mechanism for learning synaptic weights is an area of growing interest, with reward-modulated Hebbian learning (RMHL) [26, 32] as a promising candidate. RMHL, a form of three-factor Hebbian learning [17, 24], offers a potentially powerful mechanism for credit assignment without explicit backpropagation of exact gradients.
@@
Despite theoretical advances, few noise-based methods have been adapted to real-world tasks. Some attempts which have been made include application to control problems [4], for spatiotemporal pattern generation [23], and for training non-differentiable spiking neural networks training [12], and these particularly struggle with delayed reward regimes. This gap has limited the integration of noise-based approaches for effective credit assignment in temporally extended setups.
In this work, we address this gap by proposing a novel noise-driven learning mechanism for reinforcement learning (RL), referred to as noise-based reward-modulated learning (NRL), which leverages noise to approximate gradients for synaptic updates derived from reward maximization. NRL integrates key concepts from neuroscience, such as eligibility traces [19], which link past neural states to future rewards, and reward prediction error (RPE) [41, 40, 34], which drives learning toward increasingly more advantageous behaviors. We employ directional derivatives for gradient approximation, which is realized through stochastic neurons that are compatible with the inherently noisy nature of brains and other physical systems. NRL estimates the directional derivative via two forward passes: a “noisy” pass with stochastic neurons and a “clean” pass without noise. In scenarios where a noiseless pass is infeasible, as in noisy analog hardware, multiple noisy passes can be averaged to approximate the clean pass, maintaining performance. We further validate our approach on RL benchmarks with both immediate- and delayed-reward problems. Finally, we discuss the broader implications for machine learning, neuromorphic computing, and neuroscience.
@@ 2 Methods
2.1 Agent-environment setup and learning signals
We consider an agent interacting continuously with a dynamic environment, whose state at time $t$ is denoted by $S_{t}$. The agent observes $S_{t}$ and responds by selecting actions, $A_{t}$, which influence the environment. Each action or sequence of actions yields feedback in the form of positive or negative rewards $r_{t}$. This observe-act loop provides a feedback loop which can be leveraged to enable an agent to learn and adapt to the changing environment, aiming to maximize the rewards associated with its actions.
The agent’s decision-making is guided by a policy $\pi$, representing the probability of choosing an action, $a$, in the state $S_{t}$, i.e., $\pi(a\mid S_{t})$. This policy reflects the agent’s understanding of optimal actions in each state. To determine an action, the agent computes a probability distribution over the set of possible actions, modeled as a categorical distribution, where $\sum_{a}\pi(a\mid S_{t})=1$, such that
$P(A_{t}=a\mid S_{t})=\pi(a\mid S_{t})\,.$ (1)
Rewards, $r_{t}$, from the environment, received after each action or sequence of actions, are used to compute the learning signal. Internally, the agent maintains a prediction of reward, denoted by $\bar{r}_{t}$. This prediction could be formed by an arbitrary computation, but for simplicity it is modeled as a running average over recent rewards
$\bar{r}_{t}=\bar{r}_{t-1}+\lambda\left(r_{t}-\bar{r}_{t-1}\right)$ (2)
where $\lambda$ is a smoothing factor. The mismatch between the predicted reward and the actual reward, denoted by
$\delta r_{t}=r_{t}-\bar{r}_{t}$ (3)
serves as the primary learning signal. It can be interpreted as a reward prediction error (RPE) in neuroscience contexts and can be thought of as a global dopaminergic signal that modulates learning in the brain [41, 40, 34]. While conceptually similar to biological RPEs, this formulation is a simplification due to the straightforward running average used for reward prediction. Nonetheless, it effectively captures the core ideas of the proposed learning mechanism. In the Discussion, we explore how this prediction error could be refined.
@@ 2.2 Derivation of the learning rule
This section presents a complete derivation of the NRL update rule. We begin by establishing a gradient-based learning rule rooted in an optimization target that aims to maximize unexpected rewards from the environment by increasing the reward prediction error. Maximizing the RPE rather than the direct reward enables the agent to adapt dynamically to changing environments, improve long-term performance by continuously seeking rewards that exceed expectations, and maintain robustness against variations in reward structures [48, 36]. Building on this foundation, we transition to a gradient-free formulation by leveraging a directional derivative framework, where stochastic noise within the network approximates gradients. This eliminates the need for explicit backpropagation and feedback phases, enabling learning through forward passes alone. Finally, we extend our noise-based approach to handle scenarios with delayed rewards, a hallmark of real-world problems. This extension enables NRL to adapt to environments where feedback is only obtained after a series of actions.
@@ 2.2.1 Gradient-based learning rule
The agent is represented by a neural network which processes input observations and generates a probability distribution over possible actions. This neural network consists of multiple layers, each performing a linear transformation of its inputs, followed by a non-linear transformation as
$x^{l}=f(h^{l})=f(W^{l}x^{l-1})$ (4)
where $x^{l}$ represents the output of layer $l$, $h^{l}$ represents the layer pre-activation, $W^{l}$ is the weight matrix, and $f(\cdot)$ is a non-linear activation function.
In this setup, the goal is to optimize the weights $W=\{W^{1},\ldots,W^{l}\}$ across the $L$ layers in the neural network to maximize the received rewards. Our derivation begins by introducing a generic variable $\theta$, which will later be mapped to the specific neural network parameters $W$. The primary learning objective is to maximize the expected RPE, $\delta r_{t}$, which aims to achieve higher-than-anticipated rewards. For now, let us consider learning to maximise RPE at a specific timepoint $t$. Since the reward prediction is a running average of recent rewards, the learning rule seeks to outperform previously received rewards. We express this objective in terms of the policy as
$J(\theta)=\mathbb{E}_{\pi_{\theta}}\left[\delta r_{t}\right]\,.$ (5)
To optimize $J(\theta)$, we can incrementally update the weights using the gradients with respect to $\theta$ as
$\theta\leftarrow\theta+\eta\nabla_{\theta}J(\theta)$ (6)
@@
$\nabla_{\theta}J(\theta)=\mathbb{E}_{\pi_{\theta}}\left[\nabla_{\theta}\log\pi_{\theta}\left(a\mid S_{t}\right)\delta r_{t}\right]\,.$ (7)
For empirical estimation, in the case of a single sample, we approximate the gradient as
$\nabla_{\theta}J(\theta)\approx\nabla_{\theta}\log\pi_{\theta}\left(a\mid S_{t}\right)\delta r_{t}\,.$ (8)
With this approximation, we define the parameter updates as
$\theta\leftarrow\theta+\eta\;\nabla_{\theta}\log\pi_{\theta}\left(a\mid S_{t}\right)\;\delta r_{t}.$ (9)
Equation (9) resembles the REINFORCE update rule [54] but differs by using the reward prediction error (RPE), $\delta_{r}$, as the learning signal instead of cumulative rewards over full trajectories. This RPE-based approach leverages immediate feedback from rewards as they are obtained rather than requiring a full trial completion to estimate the policy gradient. It shares conceptual similarities with actor-critic methods [1], where the policy is adjusted using a temporal difference error. However, we approximate future rewards with a running average of past rewards instead of a critic network, maintaining adaptive feedback benefits with a simpler implementation.
@@ 2.2.2 Noise-based learning rule
The learning rule derived in the previous section still relies on gradient descent. To avoid using backpropagation (BP) to compute the gradients, we propose a noise-based alternative that extends Equation (9). This approach leverages gradient approximation via directional derivatives, as detailed in Appendix A, enabling a theoretically rigorous derivation of noise-driven learning.
@@
A directional derivative quantifies the rate of change of a function in a specified direction. In neural networks, we implement this concept by introducing random noise into parameters. By comparing the network’s parameters with and without this noise, we obtain an estimate of the gradient direction. To formalize this, we express the gradient term in terms of directional derivatives using Theorem 31 in Appendix A as
$\nabla p(\theta)$ $=\nabla_{\theta}\log\pi_{\theta}(a\mid S_{t})=n\mathbb{E}\left[v\nabla_{v}p(\theta)\right]$ (10)
where $v=\ifrac{\epsilon}{||\epsilon||}$ is a normalized direction vector derived from noise $\epsilon\sim\mathcal{N}(0,\sigma^{2}I_{n})$ with $n$ the number of parameters.
@@
We may expand the above via Theorem A.4, and add a difference term (parameters with noise versus without noise) to approximate the gradient by sampling under an empirical distribution
$\nabla p(\theta)=n\sum_{i=1}^{K}\left[\frac{\epsilon^{(i)}}{||\epsilon^{(i)}||^{2}}\left(p\bigl{(}\tilde{\theta}^{(i)}\bigr{)}-p\bigl{(}\theta\bigr{)}\right)\right]\,.$ (11)
@@
where $K$ denotes the number of samples and $\tilde{\theta}^{(i)}=\theta+\epsilon^{(i)}$ the noise-perturbed parameters. In practice, we consider $K=1$, analogous to single-sample updates in stochastic gradient descent. Here, $p\bigl{(}\tilde{\theta}\bigr{)}=\log\pi_{\tilde{\theta}}(a\mid S_{t})$ represents the log of the noise-perturbed output, while, $p\left(\theta\right)=\log\pi_{\theta}(a\mid S_{t})$ is the log of the noise-free output. Thus, $\rho=\log\pi_{\tilde{\theta}}(a\mid S_{t})-\log\pi_{\theta}(a\mid S_{t})$ captures the impact of the noise on the policy. Putting this together for the $K=1$ case, we obtain our noise-based learning rule
$\theta\leftarrow\theta+\eta\delta_{r}\epsilon^{*}\rho$ (12)
with $\epsilon^{*}=\ifrac{\epsilon}{||\epsilon||^{2}}$. For convenience, we absorbed the term $n$ into the learning rate $\eta$.

### v2

@@ Abstract
Biological neural systems efficiently learn from delayed rewards despite relying on noisy synaptic transmission and lacking centralized optimization mechanisms. In contrast, artificial neural networks trained with reinforcement learning typically rely on backpropagation (BP), which limits their use in resource-constrained systems or with non-differentiable components. While noise-based alternatives, like reward-modulated Hebbian learning (RMHL), provide a biologically grounded framework for credit assignment, they struggle with temporal delays and hierarchical processing —key challenges in real-world learning. In this work, we derive a novel noise-based learning rule to address these challenges. Drawing inspiration from biological neural circuits, our method uses reward prediction errors as its optimization target to generate increasingly advantageous behavior, and incorporates an eligibility trace to facilitate retrospective credit assignment. Its formulation relies on local information, aligning with biological constraints and enabling neuromorphic implementation. Experimental validation on reinforcement tasks (immediate and delayed rewards) shows our approach significantly outperforms RMHL and achieves performance comparable to BP, although with slower convergence due to its noise-driven updates. While tested on simple architectures, the results highlight the potential of noise-driven, brain-inspired learning for low-power adaptive systems, particularly in scenarios where energy efficiency and biological plausibility are a priority. These findings also offer mechanistic insights into how dopamine-like signals and synaptic stochasticity may jointly enable learning in biological networks, bridging computational models with neurobiological principles.
@@ 1 Introduction
The brain’s remarkable ability to learn and adapt to its environment is believed to arise from synaptic plasticity [4, 3, 37]. These local modifications allow neural circuits to acquire complex sensorimotor skills despite inherent challenges such as noisy neural communication [12], temporally delayed feedback/reward [23], and no centralized error-propagation mechanism to drive weight updates [48]. Understanding the mechanisms underlying this robust learning remains a central challenge in neuroscience and is increasingly inspiring novel approaches in artificial intelligence [11, 38].
When it comes to learning in artificial neural networks, backpropagation (BP) underpins modern deep learning [45]. However, its implementation in biological circuits is often seen as challenging [34]. BP requires precise, non-local error gradients, which contrasts with the locality of synaptic plasticity that depends on pre- and post-synaptic activity and neuromodulatory factors —akin to three-factor learning [21, 29]. Moreover, its deterministic nature is contrary to the inherent stochasticity of neural circuits [12, 51, 39], and its requirement for continuous, immediate error feedback contradicts biological learning, which often involves delayed feedback following a sequence of actions [33].
While some works have investigated how variants of backpropagation might be implemented in cortical microcircuits —employing predictive coding, feedback alignment, or dendritic error signals to carry nonlocal gradients [60, 59, 31]— others have proposed different plasticity rules inspired directly from neurobiology [46, 25, 63]. Among these approaches, noise‐based learning methods are especially promising. They leverage fluctuations in neural activity [19, 24, 61, 58, 16] or synaptic weights [13, 65, 7, 10], for estimating gradients. By injecting small random perturbations, these methods avoid the need for perfectly matched feedback pathways and rely only on information that biological synapses could plausibly access, aligning more closely with mechanisms observed in cortical neurons [50, 36].
@@
Given the prevalence of noise in biological neural systems [12, 51, 39], using noise as a mechanism for learning synaptic weights is an area of growing interest, with reward-modulated Hebbian learning (RMHL) [32, 40] as a promising candidate. RMHL, which is a form of three-factor Hebbian learning [21, 29], offers a potentially powerful mechanism for credit assignment without explicit backpropagation of errors. Despite theoretical advances, few noise-based methods have been adapted to real-world tasks. Some attempts which have been made include application to control problems [6], for spatiotemporal pattern generation [28], and for training non-differentiable spiking neural networks training [15], and these particularly struggle with delayed reward regimes. This gap has limited the integration of noise-based approaches for effective credit assignment in temporally extended setups.
To bridge these gaps, we propose noise-based reward-modulated learning (NRL), a biologically-inspired, gradient-free learning method compatible with delayed feedback/rewards. NRL builds on established neuroscientific pillars. It employs eligibility traces [23, 26], which tag synapses based on recent pre/post activity and noise perturbations, enabling retrospective credit assignment over behavioral timescales. It also embraces reward prediction error (RPE), that is, dopamine-like signals [48, 47, 42] that modulate these traces to reinforce actions that yield unexpected positive rewards, driving learning toward increasingly more advantageous behaviors. We derive our learning rule from first principles, employing directional derivatives to compute a local gradient estimate at each synapse, rather than relying on explicit gradients. These directional derivatives are implemented through stochastic neurons, aligning with the inherently noisy nature of neural circuits and other physical systems. NRL estimates the directional derivative with two forward passes: a “noisy” pass with stochastic neurons and a “clean” pass without noise. In scenarios where a noiseless pass is infeasible, multiple noisy passes can be averaged to approximate the clean pass, maintaining performance.
We deliberately choose an abstract modelling framework so that our experiments isolate the learning rule’s behavior without introducing additional biological complexities. We employ rate-based neurons instead of spiking models, such that the dynamics reduce to smooth activation functions and continuous outputs. We also avoid distinguishing excitatory and inhibitory populations, allowing every unit to follow the same update equation. Finally, we treat potentiation and depression as two ends of a single weight-change continuum rather than separate processes, so that each synapse simply strengthens or weakens in proportion to its eligibility trace and reinforcement signal. In this way, we trade-off biological detail for more clear insight into capabilities and limitations of our proposed noise-based learning.
We validate NRL on a suite of reinforcement-learning benchmarks spanning immediate-reward and delayed-reward tasks, demonstrating significantly superior learning efficiency when compared to RMHL (particularly when rewards were delayed by many steps), and competitive final performance compared to gradient-based baselines. Moreover, unlike RMHL, which breaks down in deeper architectures, NRL scales to multilayer networks, though converging more slowly due to its intrinsic stochasticity. Finally, we explore how these findings shed light on biological learning and point toward energy-efficient implementations on neuromorphic hardware, as well as novel, alternative, robust learning schemes for machine learning.
@@ 2 Methods
2.1 Neural system-environment interaction and reward signaling
We model a neural system within a dynamic environment, grounded in the reinforcement learning framework [54]. In this framework, the environment’s state at time $t$ is denoted by $s_{t}$. The neural system perceives $s_{t}$ and responds by initiating actions, $a_{t}$, which influence the environment. Each action or sequence of actions results in feedback in the form of positive or negative rewards $r_{t}$. This iterative process of perception, action, and outcome forms a feedback loop that enables the system to learn and adapt to the changing environment, ultimately seeking to maximize the rewards it receives.
The system’s decision-making is guided by a policy $\pi$, representing the probability of choosing an action, $a_{t}$, in a given state $s_{t}$ at time $t$, i.e., $\pi(a_{t}\mid s_{t})$. This policy captures the system’s learned strategy for choosing actions that are likely to yield favorable outcomes. To determine an action, the system computes a probability distribution across the set of possible actions, modelled as a categorical distribution $P(a_{t}\mid s_{t})\equiv\pi(a_{t}\mid s_{t})$ with $\sum_{a\in A}\pi(a\mid s_{t})=1$, where $A$ denotes the set of all possible actions.
The rewards $r_{t}$, obtained from the environment, serve as crucial learning signals. Internally, the system maintains a prediction of the expected reward, denoted by $\bar{r}_{t}$. While this prediction can be calculated through various mechanisms, we model it here for simplicity as a running average of recent rewards:
$\bar{r}_{t+1}=\bar{r}_{t}+\lambda\left(r_{t+1}-\bar{r}_{t}\right)$ (1)
where $\lambda$ is a smoothing factor that governs the influence of past rewards on the current prediction. The mismatch between the actual reward and the predicted reward, known as the reward prediction error, is then computed as
$\delta_{t}=r_{t}-\bar{r}_{t}.$ (2)
This RPE signal is a central element of our model and draws a strong parallel to the phasic activity of dopamine neurons in the brain [48, 47, 42]. Dopamine neurons are known to exhibit increased activity for unexpected rewards and decreased activity for the unexpected absence of reward [17]. Although our running average is a simplified representation of how reward expectations are formed in biological systems, it effectively captures the fundamental dynamics of the RPE and its role as a learning signal. We will explore potential refinements of this reward prediction mechanism in the Discussion section.
@@ 2.2 Derivation of the learning rule
This section presents a complete derivation of the NRL update rule. We begin by establishing a gradient-based learning rule rooted in an optimization target that aims to maximize unexpected rewards from the environment by increasing the reward prediction error. Using RPE, rather than the direct reward, enables our system to adapt dynamically to changing environments by tracking how rewards deviate from expectations, improving long-term performance by continuously seeking rewards that exceed expectations, and maintaining robustness against variations in reward structures [54, 43]. Building on this foundation, we transition to a directional derivative framework, where stochastic noise within the network approximates gradients. This eliminates the need for explicit backpropagation and feedback phases, enabling learning through forward passes alone. Finally, we extend our noise-based approach to handle scenarios with delayed rewards, a hallmark of real-world problems. This extension enables NRL to adapt to environments where feedback is only obtained after a series of actions.
@@ 2.2.1 Gradient-based learning rule
The system is modelled as a multi-layered network of interconnected units, analogous to populations of neurons, which processes input representing sensory observations and produces a probability distribution over possible actions. For simplicity, we refer to these populations as ‘layers,’ where each layer performs a transformation of its input in two stages: first, a linear transformation, followed by a non-linear transformation. This can be expressed mathematically as
$x^{l}_{t}=f(h^{l}_{t})=f(W^{l}_{t}x^{l-1}_{t})$ (3)
where $x^{l}_{t}$ represents the output activity of layer $l$, $h^{l}_{t}$ represents the layer pre-activation (the combined input to such layer), $W^{l}_{t}$ represents the matrix of synaptic weights connecting layer l1l-1 to layer $l$, at time $t$, and $f(\cdot)$ is a non-linear activation function, which introduces crucial non-linearities into the network’s computations
In this setup, the goal is to adjust the synaptic weights, $\{W^{1},\ldots,W^{L}\}$, across the $L$ layers of the network to maximize the rewards obtained by the system. Our derivation begins by introducing a general parameter $\theta$, which will later be mapped to the specific neural network parameters $W$. The primary learning objective is to maximize the RPE, $\delta_{t}$, at any particular moment in time, $t$, effectively driving the system to seek out actions that lead to higher-than-predicted rewards. The reward prediction is modeled as a running average of recent rewards and, as such, the learning rule seeks to outperform previously received rewards. Note that for a fixed reward prediction, this is equivalent to maximizing the reward itself. We express this objective in terms of a parameterized policy at time $t$ as
$J(\theta_{t})=\mathbb{E}_{\pi_{\theta_{t}}}\left[\delta_{t}\right]\,.$ (4)
To optimize $J(\theta_{t})$, we can incrementally update the weights using the gradients with respect to $\theta_{t}$ as
$\theta_{t+1}\leftarrow\theta_{t}+\eta\nabla J(\theta_{t})$ (5)
@@
$\nabla J(\theta_{t})=\mathbb{E}_{\pi_{\theta_{t}}}\left[\nabla\log\pi_{\theta_{t}}\left(a_{t}\mid s_{t}\right)\delta_{t}\right]\,.$
For empirical estimation, in the case of a single sample, we approximate the gradient as $\nabla J(\theta_{t})\approx\nabla\log\pi_{\theta_{t}}\left(a_{t}\mid s_{t}\right)\delta_{t}$. Using this approximation, we define the parameter updates as
$\theta_{t+1}\leftarrow\theta_{t}+\eta\;\nabla\log\pi_{\theta_{t}}\left(a_{t}\mid s_{t}\right)\;\delta_{t}\,.$ (6)
Equation (6) resembles the REINFORCE update rule [61] but differs by using the reward prediction error, $\delta_{t}$, as the learning signal instead of cumulative rewards over full trajectories. This RPE-based approach leverages immediate feedback from rewards as they are obtained rather than requiring a full trial completion to estimate the policy gradient. It shares conceptual similarities with actor-critic methods in reinforcement learning [1], where the policy is adjusted using a temporal difference error. However, we approximate future rewards with a running average of past rewards instead of a critic network, a less powerful but simpler implementation which maintains adaptive feedback.
@@ 2.2.2 Noise-based learning rule
The learning rule derived in the previous section still relies on gradient descent. To avoid using backpropagation to compute the gradients, we propose a noise-based alternative that extends Equation (6). This approach leverages gradient approximation via directional derivatives, enabling a theoretically rigorous derivation of noise-driven learning.
@@
A directional derivative quantifies the rate of change of a function in a specified direction. In our neural system, we implement this concept by introducing random noise into parameters. By comparing the network’s parameters with and without this noise, we obtain an estimate of the gradient direction. To formalize this, we define $g(\theta_{t})=\log\pi_{\theta_{t}}(a_{t}\mid s_{t})$ and express the gradient term in terms of directional derivatives using the theorems in Appendix A as
$\nabla g(\theta_{t})$ $=\nabla\log\pi_{\theta_{t}}(a_{t}\mid s_{t})=n\mathbb{E}\left[\bar{\epsilon}_{t}\nabla_{\bar{\epsilon}_{t}}g(\theta_{t})\right]$ (7)
[…]
@@
We may expand the above to approximate the gradient via a finite-difference and by sampling under an empirical distribution
[…]
@@
where $K$ denotes the number of samples and $\tilde{\theta}_{t}^{(i)}=\theta_{t}+\epsilon^{(i)}_{t}$ the noise-perturbed parameters (see Appendix A). In practice, we consider $K=1$, analogous to single-sample updates in stochastic gradient descent. Here, $g\bigl{(}\tilde{\theta}^{(i)}_{t}\bigr{)}=\log\pi_{\tilde{\theta}_{t}}(a_{t}\mid s_{t})$ and $g\left(\theta_{t}\right)=\log\pi_{\theta_{t}}(a_{t}\mid s_{t})$ are the log of the noise-perturbed and noise-free output, respectively. Thus, we define $\rho_{t}=\log\pi_{\tilde{\theta}_{t}}(a_{t}\mid s_{t})-\log\pi_{\theta_{t}}(a_{t}\mid s_{t})$, which captures the impact of the noise on the policy. Putting this together for the $K=1$ case, we obtain our noise-based learning rule
$\theta_{t+1}\leftarrow\theta_{t}+\eta\delta_{t}\hat{\epsilon}_{t}\rho_{t}$
[…]
2.2.3 Learning with node level noise in neural network
The noise-based learning rule derived so far is formulated for a general parameter $\theta_{t}$. Now, we apply this framework to the specific context of a neural network, where $\theta_{t}$ corresponds to the synaptic weights $W^{l}_{t}$ of layer $l$ at time $t$.
Instead of directly perturbing the weights of the network $W^{l}_{t}$, we propose extending the learning rule by introducing noise directly into the neurons (nodes) of each layer. This strategy is advantageous because it can lead to reduced variance in gradient estimation, as perturbations occur in a lower-dimensional space (the neural activations) compared to the full weight space. This approach shares similarities with node perturbation (NP) [19, 24, 61, 58, 16], which has shown benefits in reducing gradient variance and offering more localized updates.